# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ Abstract
Symmetry topological field theory (SymTFT), or topological holography, offers a unifying framework for describing quantum phases of matter and phase transitions between them. While this approach has seen remarkable success in describing gapped and gapless pure-state phases in $1+1$d, its applicability to open quantum systems remains entirely unexplored. In this work, we propose a natural extension of the SymTFT framework to mixed-state phases by introducing the symmetry taco: a bilayer topological order in $2+1$d whose folded geometry naturally encapsulates both strong and weak symmetries of the $1+1$d theory. We use this perspective to identify a series of correspondences, including a one-to-one map between intrinsically gapless SPTs (igSPTs) and certain gapped SPTs, and a mapping between igSPTs and intrinsically average SPTs (iASPTs) arising in $1+1$d mixed states. More broadly, our framework yields a classification of short-range correlated $G$-symmetric Choi states in $1+1$d, provides a route for systematically generating mixed-state SPTs via local decoherence of igSPTs, and allows us to identify a new mixed-state “anomaly”. Besides folding in mixed-state phases into the SymTFT paradigm, the symmetry taco opens new avenues for exploring dualities, anomalies, and non-equilibrium criticality in mixed-state quantum matter.
[…]
@@ I Introduction
As Anderson famously observed, “It is only slightly overstating the case to say that physics is the study of symmetry” [1]. This is especially true in the context of many-body physics, where symmetry plays a central role in organizing and classifying phases of matter. Within the Landau paradigm, phases are distinguished by distinct patterns of spontaneous symmetry breaking of a given symmetry. In contrast, symmetry protected topological (SPT) phases are quantum phases that fall outside this paradigm: they do not spontaneously break any symmetry, yet remain distinct from each other and from trivial, short-range entangled states provided a global symmetry is enforced [2, 3]. Recently, generalized notions of symmetry have led to fresh perspectives on familiar systems: for instance, topological orders (TOs)–long-range entangled phases which support fractionalized anyon excitations–do not require the imposition of any microscopic (UV) symmetries but can nevertheless be cast as phases of matter arising from the spontaneous breaking of emergent higher-form symmetries, which act on extended, rather than point-like, objects. Separately, non-local duality transformations such as the paradigmatic Kramers-Wannier duality [4] have been reinterpreted as “non-invertible” symmetries, described not by symmetry groups but rather by fusion categories [5, 6]. Such generalized notions of symmetry build on the one-to-one correspondence between a symmetry group $G$ and codimension-1 topological defects, which fuse according to group multiplication rules. From this perspective, every topological defect in a theory–which can have higher codimension and can obey non-invertible fusion rules–corresponds to a generalized global symmetry [7, 8, 9]. Such defects are expected to be described via fusion higher-categories [10, 11, 12, 13, 14, 15, 16].
@@
While such generalized notions of symmetry and the corresponding phases of matter they characterize initially appear disparate, they can in fact be unified within a single framework known as the symmetry topological field theory (SymTFT) (and its closely related cousin, topological holography[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. Within this framework, the different phases of a system with a given symmetry structure are described in terms of a topologically ordered “bulk” in one higher dimension. The SymTFT perspective has most thoroughly been studied in the context of $1+1$d systems (see Refs. [46, 47, 48, 49, 50, 51] for higher-dimensional generalizations). The central concept of this framework is that the symmetry data of a $1+1$d system with global symmetry $G$ can be encoded, via the bulk-boundary correspondence, within a $2+1$d slab with TO given by the quantum double $\mathcal{D}(G)$, with appropriately chosen boundary conditions [^1] Any theory in $d+1$ spacetime dimensions with symmetry $\mathcal{C}$ (a fusion $d$-category in general) can be recovered via its SymTFT, which is the interval compactification of a $d+2$ TO given by the Drinfeld center $\mathcal{Z}_{1}(\mathcal{C})$ with the following boundary conditions: on one end (called the “symmetry” boundary), topological boundary conditions corresponding to the canonical gapped boundary of $\mathcal{Z}_{1}(\mathcal{C})$ are enforced, while the other boundary (called the “physical” boundary) is allowed to have arbitrary non-topological boundary conditions..
@@
Different phases of the $G$-symmetric $1+1$d theory can then be identified with different choices of boundary conditions (i.e., anyon condensates) on the physical boundary of the $\mathcal{D}(G)$ TO. Crucially, properties of the symmetry operators of the $1+1$d theory–such as their anomalies–can be precisely identified with properties of the anyon operators in the SymTFT. For instance, the SymTFT of a $1+1$d $\mathbb{Z}_{2}$ symmetric system is given by the $2+1$d $\mathbb{Z}_{2}$ Toric Code, the TO of which is given by $\mathcal{D}(\mathbb{Z}_{2})=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. The different phases of the former–the symmetry breaking and symmetry preserving gapped phases, as well as symmetric gapless phases–can be identified with the different boundary anyon condensates of the latter. Moreover, the Kramers-Wannier duality of the $1+1$d theory can be seen as arising from the electromagnetic duality of the Toric Code.
@@
An appealing aspect of this framework is that it encodes all topological or kinematic properties of a $G$-symmetric theory in the corresponding higher-dimensional SymTFT while disregarding non-universal dynamical data, providing a route to making non-perturbative statements about phases and phase transitions. Consequently, this framework not only reproduces the full classification of gapped $1+1$d $G$-symmetric phases [23, 22, 24, 52, 40] but also neatly incorporates duality transformations as changes of boundary conditions, which correspond to certain anyon permutation symmetries of the bulk TO [17, 21]. More surprisingly, the SymTFT has also proven useful in classifying gapless phases of matter [36, 37, 39, 38, 42, 43, 44, 45], especially intrinsically gapless SPTs (igSPTs) [53, 54]. Despite their gapless nature, igSPTs support topological edge modes, which are determined by an emergent anomaly that arises at low energies. As we will review, igSPTs correspond to partially gapped boundaries of the bulk TO. While the SymTFT does not capture dynamical properties of igSPTs, it does describe their emergent anomaly and hence also their topological features. Given that the SymTFT can describe such a diverse array of gapless and gapped phases of matter, it is natural to ask whether we can also use this framework to establish correspondences between their classifications. This will be one of the primary themes of the present work.
@@
An important goal of this work is to provide the first step towards extending the SymTFT framework to the $1+1$d mixed-state setting; along the way, this will lead us to argue for correspondences between both gapless and gapped pure-state phases with intrinsically mixed-state phases. Our central results are illustrated in the tetraptych of Fig. 1. We first establish a one-to-one correspondence between igSPTs and certain bilayer gapped phases. We then move to the open system setting, arguing that the bilayer gapped SPTs arising from igSPTs have a natural interpretation as Choi states (i.e., vectorizations) of mixed-state SPTs. Following recent work on classifying mixed-state phases (see e.g., Refs. [74, 75]), we restrict our analysis to mixed states whose corresponding Choi states are gapped in the usual (pure state) sense. We conclude that the appropriate SymTFT for such mixed states is given by the symmetry taco (SymTaco), i.e., a bilayer TO whose two layers describe the “ket” and “bra” spaces, which are not fully independent (hence a taco and not a sandwich). This allows us to establish a correspondence between mixed-state iASPTs and pure-state igSPTs, as well as more generally provide a partial classification of all short-range correlated $1+1$d mixed states. Our perspective motivates a systematic approach for generating certain gapped SPTs via deformations of igSPT Hamiltonians and of obtaining iASPTs by decoherence of igSPTs. Finally, we also show how the symmetry taco can be further exploited to explore novel forms of gauging unique to the mixed-state setting. Before proceeding, we first summarize our main results below.
@@ I.1 Summary of Main Results
[…]
@@
To understand the utility of the SymTaco, consider the quantum double $\mathcal{D}(G)$ describing the $2+1$d bulk TO. Given this anyon theory and a condensable subalgebra of anyons $\mathcal{A}$, anyon condensation [^2] In this paper, we will only be discussing the algebraic procedure of anyon condensation and not the energetic process of tuning a Hamiltonian across a phase transition separating distinct topological orders, which is oftentimes referred to by the same moniker. refers to a formal map between two gapped phases–the original uncondensed TO $\mathcal{D}(G)$ and the condensed TO $\mathcal{D}(G)/\mathcal{A}$. This map provides a systematic approach for identifying the anyons in the former that become condensed (identified with the vacuum superselection sector), are confined, or split into distinct superselection sectors in the latter (see Ref. [76] for a discussion). The classification of condensable subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$ is well understood [77] and the condensed TO $\mathcal{D}(G)/\mathcal{A}$ is generally described by a twisted quantum double $\mathcal{D}_{\omega}(G/K)$ specified by a normal subgroup $K\lhd G$ and $[\alpha]\in H^{3}(G/K,U(1))$ a 3-cocycle class [78]. Besides the bulk picture, anyon condensation can also be understood holographically, i.e., in terms of states defined on a $1+1$d boundary; we now outline three distinct physical pictures that follow from this holographic dictionary.
@@
Suppose we perform anyon condensation specified by the condensable subalgebra $\mathcal{A}$, but only in some connected subregion of the $\mathcal{D}(G)$ TO–this leads to the subregion containing $\mathcal{D}(G)/\mathcal{A}$ TO, which is separated from the original TO by a pair of gapped domain walls (each specified by a module category [79]), as shown in Fig. 1. We can then pick these domain walls in a canonical manner such that anyons from $\mathcal{D}(G)/\mathcal{A}$ can pass through a domain wall into the $\mathcal{D}(G)$ TO without leaving any nontrivial excitations that are confined on the domain wall [38]. Now, we fold this $2+1$d system with a subregion containing the condensed anyon theory into a taco, as illustrated in Fig. 1. In accordance with the folding trick, the bulk TO is now doubled, i.e., it is given by $\mathcal{D}(G^{\prime})$ (with $G^{\prime}=G\times G$) [^3] This follows from the fact that $\mathcal{Z}(G\times G)\cong\mathcal{Z}(G)\boxtimes\mathcal{Z}(G)$ [77, 80].. We will refer to this bulk TO folded along the condensed subregion as the SymTaco.
@@
The first way to interpret the condensed subregion via this SymTaco  is as follows: after folding, we obtain a gapped boundary between the folded $\mathcal{D}(G^{\prime})$ TO and vacuum, with the particular choice of gapped boundary determined by the condensable algebra $\mathcal{A}$. Thus, distinct allowed choices of gapped boundaries between the folded TO and vacuum correspond to distinct condensable subalgebras $\mathcal{A}$. On the other hand, gapped boundaries between $\mathcal{D}(G^{\prime})$ TO (not necessarily obtained via folding) and vacuum are classified via $1+1$d gapped phases with symmetry $G^{\prime}$, i.e., in terms of a subgroup $H\subset G^{\prime}$, which characterizes symmetry breaking, and a cohomology class $[\omega]\in H^{2}(H,U(1))$, which characterizes SPT order  [80]. Imposing the constraint that the bulk TO is obtained via folding, only a subset of all possible gapped boundaries between $\mathcal{D}(G^{\prime})$ and vacuum can be obtained, corresponding to a subset of all $G^{\prime}$-symmetric $1+1$d gapped phases. We characterize this subset of phases precisely in Sec. IV, and show how it relates to the classification of condensable subalgebras obtained in Ref. [77]. We remark that this correspondence between condensable subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$ and $1+1$d $G^{\prime}$ symmetric gapped phases was previously obtained in Ref. [81] using tensor network techniques.
@@
A second perspective relating anyon condensation to $G$-symmetric $1+1$d phases appears when we now break the SymTaco  in half, i.e., cut the system with the condensed subregion in half, as shown in Fig. 1. Now, we obtain a boundary between $\mathcal{D}(G)$ TO and vacuum, but this boundary is generically gapless because the algebra that we condensed is non-Lagrangian. Physically, this gaplessness can be enforced if the bulk $1$-form symmetry generated by the uncondensed anyons acts in an anomalous manner on the boundary, forbidding a trivial symmetric gapped boundary theory. In other words, the boundary theory must either break this symmetry spontaneously or it must be gapless. As discussed in Refs. [36, 37], the gapless boundary states can be interpreted as gapless SPT phases and hence, the SymTaco  also encodes the fact that the classification of $1+1$d gapless SPT phases protected by $G$ symmetry is in one-to-one correspondence with the condensable subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$ and, following the preceding paragraph, with a subclass of $1+1$d gapped phases with $G^{\prime}$ symmetry. Indeed, we show how to directly translate between the algebraic data specifying a $1+1$d gapless phase and that of the corresponding gapped phase, and discuss the physical meaning of this correspondence.
@@
Finally, we will put these results together to show that the folded boundary of the SymTaco  also describes mixed-state SPT phases, providing a third manner in which anyon condensation in the bulk theory encodes the data of $G$-symmetric $1+1$d phases and extending the SymTFT paradigm to the mixed-state setting. We obtain this novel correspondence by making a precise identification between the constraints characterizing the folded boundary and the constraints characterizing Choi states within a doubled Hilbert space. Alternatively, we can also view the folded bulk TO as a mixed-state topological order, providing a SymTFT picture for the boundary mixed state.
@@
Thus, the holographic correspondence as encoded in the SymTaco  allows us to interpret anyon condensation in $\mathcal{D}(G)$ TOs in terms of $1+1$d (i) gapped $G\times G$-symmetric phases, (ii) gapless $G$-symmetric phases, or (iii) $G$-symmetric mixed-state phases. This leads to a nontrivial correspondence between these three kinds of $1+1$d phases; for instance, it is known that $G$-symmetric gapless SPTs and mixed-state SPTs can exhibit symmetry fractionalization patterns that are impossible in $G$-symmetric gapped SPT phases [54]. Nevertheless, our correspondence provides a way of relating these phases: given a gapless/mixed-state SPT with symmetry $G$, we can determine the corresponding condensable subalgebra $\mathcal{A}$ in the bulk SymTaco, which then allows us to determine the corresponding gapped $G\times G$-symmetric phase. In this way, we can establish equivalences between the tetraptych shown in Fig. 1, which summarizes our main results.
@@
The balance of this paper is organized as follows: we review the SymTFT construction for gapped $1+1$d phases in Sec. II, followed by a review of partial anyon condensation and its relation to gapless $1+1$d SPT phases in Sec. III. In both of these review Sections, we introduce the requisite formalism and discuss explicit examples which will inform our discussion in the following Sections. We introduce the SymTaco  in Sec. IV and use it to establish our first non-trivial correspondence between $G$-symmetric $1+1$d igSPTs and a subset of gapped $G\times G$-symmetric $1+1$d phases, which can be understood as folded boundaries of a $2+1$d $\mathcal{D}(G\times G)$ TO. Here, we show how to relate the symmetry fractionalization data characterizing these seemingly distinct phases, and use this to identify the Hamiltonian deformations that take a stack of igSPTs to a folded SPT. We develop the SymTaco perspective for $1+1$d $G$-symmetric mixed states in Section V, where we show how the strong (exact) and weak (average) symmetries inherent to mixed states are encoded in the SymTFT for the Choi state. Crucially, we identify the positivity and Hermiticity constraints on density matrices as being precisely the constraints characterizing folded boundaries of the SymTaco. This in turn allows us to relate igSPTs and iASPTs, and to provide a classification for the latter (under certain physically reasonable assumptions that we clarify). Finally, in Sec. VI we discuss some applications of the SymTaco to mixed states, which include the identification of a new mixed-state “anomaly,” a description of mixed-state SPTs (for Abelian $G$) as symmetric Pauli subsystem codes, and a method for generating intrinsically average SPTs from igSPTs. We conclude in Sec. VII with a discussion of open questions and future directions.

### v2

@@ Abstract
Symmetry topological field theory (SymTFT), or topological holography, offers a unifying framework for describing quantum phases of matter and phase transitions between them. While this approach has seen remarkable success in describing gapped and gapless pure-state phases in 1+11+1d, its applicability to open quantum systems remains entirely unexplored. In this work, we propose a natural extension of the SymTFT framework to mixed-state phases by introducing the symmetry taco: a bilayer topological order in 2+12+1d whose folded geometry naturally encapsulates both strong and weak symmetries of the 1+11+1d theory. We use this perspective to identify a series of correspondences, including a one-to-one map between intrinsically gapless SPTs (igSPTs) and certain gapped SPTs, and a mapping between igSPTs and intrinsically average SPTs (iASPTs) arising in 1+11+1d mixed states. More broadly, our framework yields a classification of short-range correlated $G$-symmetric Choi states in 1+11+1d, provides a route for systematically generating mixed-state SPTs via local decoherence of igSPTs, and allows us to identify a new mixed-state “anomaly”. Besides folding in mixed-state phases into the SymTFT paradigm, the symmetry taco opens new avenues for exploring dualities, anomalies, and non-equilibrium criticality in mixed-state quantum matter.
Figure 1: The Symmetry Taco (symRefer to caption) framework: (a) starting with a 2+12+1d topological order described by the quantum double $\mathcal{D}(G)$, condense a subset of anyons–specified by a condensable subalgebra $\mathcal{A}$–in a subregion to obtain a new topological order $\mathcal{D}(G)/\mathcal{A}$. (b) Folding this theory along the mirror axis (dashed black line) produces the SymTaco. This setup describes gapped boundaries from the folded $\mathcal{D}(G)\times\overline{\mathcal{D}(G)}\cong\mathcal{D}(G\times G)$ topological order to vacuum. Such folded boundaries characterize a subset of 1+11+1d gapped phases with $G\times G$ symmetry. (c) Breaking the SymTaco  in half along the dashed black line produces a gapless boundary from the $\mathcal{D}(G)$ topological order to vacuum. Such boundaries characterize $G$-symmetric intrinsically gapless SPT phases. (d) Constraints imposed by folding along the dashed black line precisely correspond to the positivity and Hermiticity constraints on Choi states of density matrices. In the Choi space, folding corresponds to maximal decoherence which couples the two layers, represented here via the bold red lines. The SymTaco  hence provides the SymTFT for 1+11+1d $G$-symmetric mixed-state SPTs. In all cases, the symmetry boundary (from $\mathcal{D}(G)$ to vacuum) is given by the canonical charge condensed boundary. See the summary of results I.1 for details.
@@ I Introduction
As Anderson famously observed, “It is only slightly overstating the case to say that physics is the study of symmetry” [1]. This is especially true in the context of many-body physics, where symmetry plays a central role in organizing and classifying phases of matter. Within the Landau paradigm, phases are distinguished by distinct patterns of spontaneous symmetry breaking of a given symmetry. In contrast, symmetry protected topological (SPT) phases are quantum phases that fall outside this paradigm: they do not spontaneously break any symmetry, yet remain distinct from each other and from trivial, short-range entangled states provided a global symmetry is enforced [2, 3]. Recently, generalized notions of symmetry have led to fresh perspectives on familiar systems: for instance, topological orders (TOs)–long-range entangled phases which support fractionalized anyon excitations–do not require the imposition of any microscopic (UV) symmetries but can nevertheless be cast as phases of matter arising from the spontaneous breaking of emergent higher-form symmetries, which act on extended, rather than point-like, objects. Separately, non-local duality transformations such as the paradigmatic Kramers-Wannier duality [4] have been reinterpreted as “non-invertible” symmetries, described not by symmetry groups but rather by fusion categories [5, 6]. Such generalized notions of symmetry build on the one-to-one correspondence between a symmetry group $G$ and codimension-1 topological defects, which fuse according to group multiplication rules. From this perspective, every topological defect in a theory–which can have higher codimension and can obey non-invertible fusion rules–corresponds to a generalized global symmetry [7, 8, 9, 10]. Such defects are expected to be described via fusion higher-categories [11, 12, 13, 14, 15, 16, 17].
@@
While such generalized notions of symmetry and the corresponding phases of matter they characterize initially appear disparate, they can in fact be unified within a single framework known as the symmetry topological field theory (SymTFT) (and its closely related cousin, topological holography) [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Within this framework, the different phases of a system with a given symmetry structure are described in terms of a topologically ordered “bulk” in one higher dimension. The SymTFT perspective has most thoroughly been studied in the context of 1+11+1d systems (see Refs. [49, 50, 51, 52, 53, 54] for higher-dimensional generalizations). The central concept of this framework is that the symmetry data of a 1+11+1d system with global symmetry $G$ can be encoded, via the bulk-boundary correspondence, within a 2+12+1d slab with TO given by the quantum double $\mathcal{D}(G)$, with appropriately chosen boundary conditions [^1] Any theory in $d+1$ spacetime dimensions with symmetry $\mathcal{C}$ (a fusion $d$-category in general) can be recovered via its SymTFT, which is the interval compactification of a $d+2$ TO given by the Drinfeld center $\mathcal{Z}_{1}(\mathcal{C})$ with the following boundary conditions: on one end (called the “symmetry” boundary), topological boundary conditions corresponding to the canonical gapped boundary of $\mathcal{Z}_{1}(\mathcal{C})$ are enforced, while the other boundary (called the “physical” boundary) is allowed to have arbitrary non-topological boundary conditions..
@@
Different phases of the $G$-symmetric 1+11+1d theory can then be identified with different choices of boundary conditions (i.e., anyon condensates) on the physical boundary of the $\mathcal{D}(G)$ TO. Crucially, properties of the symmetry operators of the 1+11+1d theory–such as their anomalies–can be precisely identified with properties of the anyon operators in the SymTFT. For instance, the SymTFT of a 1+11+1d $\mathbb{Z}_{2}$ symmetric system is given by the 2+12+1d $\mathbb{Z}_{2}$ Toric Code, the TO of which is given by $\mathcal{D}(\mathbb{Z}_{2})=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. The different phases of the former–the symmetry breaking and symmetry preserving gapped phases, as well as symmetric gapless phases–can be identified with the different boundary anyon condensates of the latter. Moreover, the Kramers-Wannier duality of the 1+11+1d theory can be seen as arising from the electromagnetic duality of the Toric Code.
@@
An appealing aspect of this framework is that it encodes all topological or kinematic properties of a $G$-symmetric theory in the corresponding higher-dimensional SymTFT while disregarding non-universal dynamical data, providing a route to making non-perturbative statements about phases and phase transitions. Consequently, this framework not only reproduces the full classification of gapped 1+11+1d $G$-symmetric phases [24, 23, 25, 55, 41] but also neatly incorporates duality transformations as changes of boundary conditions, which correspond to certain anyon permutation symmetries of the bulk TO [18, 22]. More surprisingly, the SymTFT has also proven useful in classifying gapless phases of matter [37, 38, 40, 39, 43, 44, 45, 46], especially intrinsically gapless SPTs (igSPTs) [56, 57]. Despite their gapless nature, igSPTs support topological edge modes, which are determined by an emergent anomaly that arises at low energies. As we will review, igSPTs correspond to partially gapped boundaries of the bulk TO. While the SymTFT does not capture dynamical properties of igSPTs, it does describe their emergent anomaly and hence also their topological features. Given that the SymTFT can describe such a diverse array of gapless and gapped phases of matter, it is natural to ask whether we can also use this framework to establish correspondences between their classifications. This will be one of the primary themes of the present work.
@@
An important goal of this work is to provide the first step towards extending the SymTFT framework to the 1+11+1d mixed-state setting [^2] See also Refs. [77, 78] which discuss distinct holographic perspectives for mixed-states.; along the way, this will lead us to argue for correspondences between both gapless and gapped pure-state phases with intrinsically mixed-state phases. Our central results are illustrated in the tetraptych of Fig. 1. We first establish a one-to-one correspondence between igSPTs and certain bilayer gapped phases. We then move to the open system setting, arguing that the bilayer gapped SPTs arising from igSPTs have a natural interpretation as Choi states (i.e., vectorizations) of mixed-state SPTs. Following recent work on classifying mixed-state phases (see e.g., Refs. [79, 80]), we restrict our analysis to mixed states whose corresponding Choi states are gapped in the usual (pure state) sense. We conclude that the appropriate SymTFT for such mixed states is given by the symmetry taco (SymTaco), i.e., a bilayer TO whose two layers describe the “ket” and “bra” spaces, which are not fully independent (hence a taco and not a sandwich). This allows us to establish a correspondence between mixed-state iASPTs and pure-state igSPTs, as well as more generally provide a partial classification of all short-range correlated 1+11+1d mixed states. Our perspective motivates a systematic approach for generating certain gapped SPTs via deformations of igSPT Hamiltonians and of obtaining iASPTs by decoherence of igSPTs. Finally, we also show how the symmetry taco can be further exploited to explore novel forms of gauging unique to the mixed-state setting. Before proceeding, we first summarize our main results below.
@@ I.1 Summary of Main Results
The principal character in this work is a folded 2+12+1d bulk topological order, whose anyon theory is specified by representations of the quantum double $\mathcal{D}(G)$ (for $G$ a finite group) and whose boundary conditions we will specify later. We will show that this setup, which we dub the symmetry taco (sym[Uncaptioned image]), naturally allows us to utilize the theory of anyon condensation in the 2+12+1d TO to understand and relate gapped, gapless, and mixed-state SPTs in 1+11+1d. Note that while our focus is $\mathcal{D}(G)$ TOs here, the SymTaco  can be defined more generally for a bulk TO whose anyon theory is described by a unitary modular tensor category $\mathcal{C}$ (see Sec. IV).
@@
To understand the utility of the SymTaco, consider the quantum double $\mathcal{D}(G)$ describing the 2+12+1d bulk TO. Given this anyon theory and a condensable subalgebra of anyons $\mathcal{A}$, anyon condensation [^3] In this paper, we will only be discussing the algebraic procedure of anyon condensation and not the energetic process of tuning a Hamiltonian across a phase transition separating distinct topological orders, which is oftentimes referred to by the same moniker. refers to a formal map between two gapped phases–the original uncondensed TO $\mathcal{D}(G)$ and the condensed TO $\mathcal{D}(G)/\mathcal{A}$. This map provides a systematic approach for identifying the anyons in the former that become condensed (identified with the vacuum superselection sector), are confined, or split into distinct superselection sectors in the latter (see Ref. [81] for a discussion). The classification of condensable subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$ is well understood [82] and the condensed TO $\mathcal{D}(G)/\mathcal{A}$ is generally described by a twisted quantum double $\mathcal{D}_{\omega}(G/K)$ specified by a normal subgroup $K\lhd G$ and $\alpha]\in H^{3}(G/K,U(1))$ a 3-cocycle class [83]. Besides the bulk picture, anyon condensation can also be understood holographically, i.e., in terms of states defined on a 1+11+1d boundary; we now outline three distinct physical pictures that follow from this holographic dictionary.
@@
Suppose we perform anyon condensation specified by the condensable subalgebra $\mathcal{A}$, but only in some connected subregion of the $\mathcal{D}(G)$ TO–this leads to the subregion containing $\mathcal{D}(G)/\mathcal{A}$ TO, which is separated from the original TO by a pair of gapped domain walls (each specified by a module category [84]), as shown in Fig. 1. We can then pick these domain walls in a canonical manner such that anyons from $\mathcal{D}(G)/\mathcal{A}$ can pass through a domain wall into the $\mathcal{D}(G)$ TO without leaving any nontrivial excitations that are confined on the domain wall [39]. Now, we fold this 2+12+1d system with a subregion containing the condensed anyon theory into a taco, as illustrated in Fig. 1. In accordance with the folding trick, the bulk TO is now doubled, i.e., it is given by $\mathcal{D}(G^{\prime})$ (with $G^{\prime}=G\times G$) [^4] This follows from the fact that $\mathcal{Z}(G\times G)\cong\mathcal{Z}(G)\boxtimes\mathcal{Z}(G)$ [82, 85].. We will refer to this bulk TO folded along the condensed subregion as the SymTaco.
@@
The first way to interpret the condensed subregion via this SymTaco  is as follows: after folding, we obtain a gapped boundary between the folded $\mathcal{D}(G^{\prime})$ TO and vacuum, with the particular choice of gapped boundary determined by the condensable algebra $\mathcal{A}$. Thus, distinct allowed choices of gapped boundaries between the folded TO and vacuum correspond to distinct condensable subalgebras $\mathcal{A}$. On the other hand, gapped boundaries between $\mathcal{D}(G^{\prime})$ TO (not necessarily obtained via folding) and vacuum are classified via 1+11+1d gapped phases with symmetry $G^{\prime}$, i.e., in terms of a subgroup $H\subset G^{\prime}$, which characterizes symmetry breaking, and a cohomology class $\omega]\in H^{2}(H,U(1))$, which characterizes SPT order [85]. Imposing the constraint that the bulk TO is obtained via folding, only a subset of all possible gapped boundaries between $\mathcal{D}(G^{\prime})$ and vacuum can be obtained, corresponding to a subset of all $G^{\prime}$-symmetric 1+11+1d gapped phases. We characterize this subset of phases precisely in Sec. IV, and show how it relates to the classification of condensable subalgebras obtained in Ref. [82]. We remark that this correspondence between condensable subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$ and 1+11+1d $G^{\prime}$ symmetric gapped phases was previously obtained in Ref. [86] using tensor network techniques.
@@
A second perspective relating anyon condensation to $G$-symmetric 1+11+1d phases appears when we now break the SymTaco  in half, i.e., cut the system with the condensed subregion in half, as shown in Fig. 1. Now, we obtain a boundary between $\mathcal{D}(G)$ TO and vacuum, but this boundary is generically gapless because the algebra that we condensed is non-Lagrangian. Physically, this gaplessness can be enforced if the bulk $1$-form symmetry generated by the uncondensed anyons acts in an anomalous manner on the boundary, forbidding a trivial symmetric gapped boundary theory. In other words, the boundary theory must either break this symmetry spontaneously or it must be gapless. As discussed in Refs. [37, 38], the gapless boundary states can be interpreted as gapless SPT phases and hence, the SymTaco  also encodes the fact that the classification of 1+11+1d gapless SPT phases protected by $G$ symmetry is in one-to-one correspondence with the condensable subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$ and, following the preceding paragraph, with a subclass of 1+11+1d gapped phases with $G^{\prime}$ symmetry. Indeed, we show how to directly translate between the algebraic data specifying a 1+11+1d gapless phase and that of the corresponding gapped phase, and discuss the physical meaning of this correspondence.
@@
Finally, we will put these results together to show that the folded boundary of the SymTaco  also describes mixed-state SPT phases, providing a third manner in which anyon condensation in the bulk theory encodes the data of $G$-symmetric 1+11+1d phases and extending the SymTFT paradigm to the mixed-state setting. We obtain this novel correspondence by making a precise identification between the constraints characterizing the folded boundary and the constraints characterizing Choi states within a doubled Hilbert space. Alternatively, we can also view the folded bulk TO as a mixed-state topological order, providing a SymTFT picture for the boundary mixed state.
@@
Thus, the holographic correspondence as encoded in the SymTaco  allows us to interpret anyon condensation in $\mathcal{D}(G)$ TOs in terms of 1+11+1d (i) gapped $G\times G$-symmetric phases, (ii) gapless $G$-symmetric phases, or (iii) $G$-symmetric mixed-state phases. This leads to a nontrivial correspondence between these three kinds of 1+11+1d phases; for instance, it is known that $G$-symmetric gapless SPTs and mixed-state SPTs can exhibit symmetry fractionalization patterns that are impossible in $G$-symmetric gapped SPT phases [57]. Nevertheless, our correspondence provides a way of relating these phases: given a gapless/mixed-state SPT with symmetry $G$, we can determine the corresponding condensable subalgebra $\mathcal{A}$ in the bulk SymTaco, which then allows us to determine the corresponding gapped $G\times G$-symmetric phase. In this way, we can establish equivalences between the tetraptych shown in Fig. 1, which summarizes our main results.
@@
The balance of this paper is organized as follows: we review the SymTFT construction for gapped 1+11+1d phases in Sec. II, followed by a review of partial anyon condensation and its relation to gapless 1+11+1d SPT phases in Sec. III. In both of these review Sections, we introduce the requisite formalism and discuss explicit examples which will inform our discussion in the following Sections. We introduce the SymTaco  in Sec. IV and use it to establish our first non-trivial correspondence between $G$-symmetric 1+11+1d igSPTs and a subset of gapped $G\times G$-symmetric 1+11+1d phases, which can be understood as folded boundaries of a 2+12+1d $\mathcal{D}(G\times G)$ TO. Here, we show how to relate the symmetry fractionalization data characterizing these seemingly distinct phases, and use this to identify the Hamiltonian deformations that take a stack of igSPTs to a folded SPT. We develop the SymTaco perspective for 1+11+1d $G$-symmetric mixed states in Section V, where we show how the strong (exact) and weak (average) symmetries inherent to mixed states are encoded in the SymTFT for the Choi state. Crucially, we identify the positivity and Hermiticity constraints on density matrices as being precisely the constraints characterizing folded boundaries of the SymTaco. This in turn allows us to relate igSPTs and iASPTs, and to provide a classification for the latter (under certain physically reasonable assumptions that we clarify). Finally, in Sec. VI we discuss some applications of the SymTaco to mixed states, which include the identification of a new mixed-state “anomaly,” a description of mixed-state SPTs (for Abelian $G$) as symmetric Pauli subsystem codes, and a method for generating intrinsically average SPTs from igSPTs. We conclude in Sec. VII with a discussion of open questions and future directions. Appendix A contains a brief discussion of the various notations used throughout the text.